01-02 March 2014

Measurements, troubleshooting, and clouds

The weekend began with the Ridge Lab team shovelling out the road to PEARL between the Canadian Forces' Upper Paradise station and the lab Saturday morning. It was -41°C in Eureka and only -22°C at PEARL (-36°C with the windchill). It's hard to say exactly what the temperature was where the team was shovelling, but it felt cold! On the positive side, the visibility was 21 km, which gave the team a great view of the nearby fjord. Clouds moved in during the afternoon, however. Sunday was cloudy and surprisingly warm -26°C in Eureka and only -16°C outside the Ridge Lab.

In the PEARL Ridge Lab's IR lab, Joseph and Dan spent Saturday morning troubleshooting the suntracker. First, they resolved the issue with the computer's time. (The computer had spontaneously decided it was New Year's Eve, 1969, shortly before we left the lab on Friday.) With that issue resolved, they turned their attention to the fact that they could no longer control the dome and the tracker. With Jonathan on the phone, they went through possible problems. After a couple of fixes, the tracker was working very well on active mode. Joseph and Dan were very thankful to have Jonathan on the phone to help them troubleshoot the suntracker.

After the suntracker was up and running again, and before clouds rolled in mid-afternoon, Joseph and Dan took 30 measurements with the Bruker. Dan took 12 measurements with PARIS. With everything running well, they anticipate acquiring full days of measurements from here on out - so long as clouds aren't in the way.

The sky was covered in clouds on Sunday, so no measurements were taken with the Bruker or PARIS. Joseph ran HCl cell tests while Dan processed PARIS measurements from the last couple of days.

Over in the UV-Vis Lab, PEARL-GBS is working well. Sophie continued to work on the alignment of light going into the UT-GBS's CCD. The alignment of the fiber's slit with the spectrometer's slit seemed to be correct, but the spectra were not satisfying on the edges of the CCD. Xiaoyi advised her to move the lamp further away from the telescope to let it light the entire CCD chip. On Sunday, Sophie refocused the light in the CCD of UT-GBS following Xiaoyi's advice. She started to do resolution tests to determine if her re-alignment was correct or not. Sophie recorded the last test in the very last minute before leaving the Ridge Lab and will see tomorrow if it worked.

When Tom and Zahra arrived at the UV-Vis lab Saturday morning, they discovered that the quartz windows that we used to cover the instrument apertures to avoid snow getting into them had frosted up overnight. They will need to clean them every morning when they get to the lab. The instruments go into sleep mode overnight; during sleep mode they point themselves downwards to avoid snow covering the apertures. Before sunrise the instruments went into zenith measurement mode. They are operating nominally and measurements are being taken constantly. In zenith mode both instruments point straight up. This gives Tom and Zahra the opportunity to align them together and with the suntracker. This was done by placing a precise leveler on the instrument aperture and moving the instruments in elevation until they are both level. After aligning the instruments their screws were both tightened and measurements were resumed.

Saturday night, Tom, Paul and Zahra disconnected the Double Brewer (number 192) from its stand at the Eureka Weather Station and loaded it on the CANDAC truck. It was transported to the Ridge Lab as part of the team's Sunday commute. Once at the Ridge Lab on Sunday, Zahra scraped off the frost from the SPS and MAESTRO apertures. There wasn't much direct sunlight Sunday so SPS and MAESTRO made zenith sky measurements.

Brewer #192 was brought into the UV-Vis Lab, set up, and tested. Tom made sure that the Brewer was operational and that the instrument and tracker software was working. Tom and Zahra plan to move it to the Ridge Lab's roof tomorrow so it can start measurements.

Down at 0PAL, the CRL had a very productive weekend. The lidar took routine measurements throughout Friday/Saturday night in all UV and Vis channels ñ for a total of 15 hours. Sham and Emily spent Saturday morning preparing lamp calibrations. Over lunch, they ran dark count measurements. Saturday afternoon was spent experimenting with various ways to optimize depolarization calibrations. The lidar performed routine measurements overnight while Sham monitored the system "remotely" from the weather station. By brunch on Sunday another 15-hours of measurements in all channels had been taken.

On Sunday, Emily and Sham set up a polarization experiment by installing a large polarizer inside the instrument. Calibrations in four configurations were made with this optic. In the evening, Sham practiced her new CRL skills by starting the lidar for measurements overnight in all UV and Visible channels. Time was also spent constructing a support for yet another polarizing optic for depolarization calibration. They anticipate installing this optic on Monday.

ACE ozonesonde flights have now begun. The first balloon was launched on Friday at 6:16 PM (local time). It reached an altitude of 32 972 meters (6.8 hPa). Saturday's ozonesonde flight occurred at 6:16 PM (local time), and reached an altitude of 30 462 meters (9.2 hPa). On Sunday, it wasn't possible to launch a Raven balloon because of winds. The ozonesonde flight was launched at 6:21 PM (local time) but failed early, reaching an altitude of only 9211 meters (274 hPa).

Saturday night involved relaxation after a busy week of travel and setting up the instruments. Most of the team played a few rounds of Paul's "Zombie dice" ñ a game where you attempt to collect the most brains ñ somewhat fitting for a group of scientists. Sunday night was the traditional "Oscar Night". The team members guessed ahead of time who will win each award, even though most of the team had never heard of the majority of the movies up for honours.

Join us in wishing for clear skies this week!

- Dan
Back to Daily Reports
Back to Home